A substance of mass m1 at a temperature of 300 K is supplied with 1000 J of energy such that its temperature increases to 350 K. When the same substance of mass m2 is supplied with 800J of energy, its temperature changes from 275 K to 305 K.


Determine the ratio of m1 to m2

1
Expert's answer
2021-06-14T14:55:23-0400

"(1)"

Case first

"Q_1=1000J"

"T_1=300K"

"T_2=350K"

Mass ="m_1"


(2)Case second

"Q'_2=800J"

"T'_1=275K"

"T'_2=305K"

Mass ="m_2"

We know that

"Q=mc_v\u2206T"

Now

"Q_1=m_1c_v(T_2-T_1)\rightarrow(1)"

"Q'_2=m_2c_v(T'_2-T'_1)\rightarrow(2)"

equation (2) devided by equation (1)

"\frac{m_2}{m_1}=\frac{Q'_2}{Q_1}\times\frac{T_2-T_1}{T'_2-T'_1}"

Put value


"\frac{m_2}{m_1}=\frac{800}{1000}\times\frac{350-300}{305-275}=\frac{8}{10}\times\frac{5}{3} =\frac{4}{3}"

"\frac{m_2}{m_1}=\frac{4}{3}"